OpenStudy (anonymous):

Integrate e^3x-4/e^2x

OpenStudy (anonymous):

Let u=e^x

OpenStudy (anonymous):

\[\Large \int\limits_{}^{}e^{x} dx=e^{x}+c\]

OpenStudy (anonymous):

Also \[\Large \int\limits\limits_{}^{}e^{ax} ~dx=\frac{1}{a}e^{ax}+c\]

OpenStudy (anonymous):

In this case separate the two: \[\Large e^{3x}\] And the \[\Large -\frac{4}{e^{2x}}\]=\[\Large -4e^{-2x}\]

OpenStudy (anonymous):

I did that and then set u=e^x but I don't know where to go from there.

OpenStudy (anonymous):

I don't advise doing that. Set u=3x in the first one

OpenStudy (anonymous):

Okay I did that and now what do I do?

OpenStudy (anonymous):

\[\Large \int\limits_{}^{}e^{3x}-\frac{4}{e^{2x}}~~ dx=\int\limits_{}^{}e^{3x}~~ dx-4\int\limits_{}^{}e^{-2x}~~dx\]

OpenStudy (anonymous):

Solve for dx when you take the differential of both sides

OpenStudy (anonymous):

I set u=-2x and got -1/2du=dx. I don't know where to put the -1/2 though.

OpenStudy (anonymous):

Plug it in for the dx

OpenStudy (anonymous):

I get e^x-4e^u

OpenStudy (anonymous):

\[\Large -4\int\limits\limits_{}^{}e^{-2x}~dx=~-4\int\limits\limits_{}^{}e^{u} \times \frac{ 1 }{ 2 }du\]

OpenStudy (anonymous):

Simplifying, what do you get? We're simplifying the second equation first

OpenStudy (anonymous):

-2e^udu?

OpenStudy (anonymous):

There you go

OpenStudy (anonymous):

OK and now do the integral based on that rule \[\Large \int\limits\limits_{}^{}e^{x} dx=e^{x}+C\]

OpenStudy (anonymous):

-2e^-2x+c would be the answer.

OpenStudy (anonymous):

Correct, now we can plug that into \[\Large \int\limits\limits_{}^{}e^{3x}-\frac{4}{e^{2x}}~~ dx=\int\limits\limits_{}^{}e^{3x}~~ dx-4\int\limits\limits_{}^{}e^{-2x}~~dx\]

OpenStudy (anonymous):

\[\Large \int\limits\limits\limits_{}^{}e^{3x}-\frac{4}{e^{2x}}~~ dx=\int\limits\limits\limits_{}^{}e^{3x}~~ dx+2e^{-2x}+C\]

OpenStudy (anonymous):

So is the final answer -1/2e^3x+2e^-2x+c?

OpenStudy (anonymous):

There was a little mistake. u= -2x -1/2 du= dx

OpenStudy (anonymous):

For the second one, you have to set u= 3x and do the same operation

OpenStudy (anonymous):

@camila1315 are you still there/